Sum of n terms of any A.P is n^2+2n . 
Find the first term and common difference

Grade:12th pass

2 Answers

Arun
5 years ago
 
Dear student
 
1st term can be written as T1 = S1 = 1 + 2 = 3
 
now
S2 = T1 + T2 = 8
 
hence T2 = 8- 3 = 5
 
hence common difference = T2 – T1 = 5- 3 = 2
AKSHANSH PAREEK
5 years ago

Sn = n² + 2n;

S(n-1) = (n-1)² + (n-1) = n2-n

Tn = Sn - S(n-1) = 3n

Thus, d = Tn - T(n - 1), where Tn-1 = 3n-1

Hence, d = 3n-3n+1

=1     and first term isTn=a+(n-1)1

                                      3n =a+n-1 

                                       a=2n+1

please approve